OpenStudy (anonymous):

if 30 ml of 1.50 M cacl2 is added to 15.0 ml of 0.100 M Agno3, what is the mass in grams of agcl precipitate

OpenStudy (joannablackwelder):

Concentration*volume = moles as long as the units cancel out :)

OpenStudy (joannablackwelder):

Using this info, can you calculate the moles of each reactant?

OpenStudy (anonymous):

mole of Cl- =\[0.03 \times 1.5 \times 2=0.09\] mole of Ag+ =\[0.015 \times 0.1=0.0015\] Ag+ is limiting reactant so we have 0.0015 mole AgCl. \[mole \times molar mass = mass of perticipate\] \[0.0015 \times 143.31=0.215 gr\]

OpenStudy (anonymous):

CaCl2 + 2 AgNO3 → 2 AgCl + Ca(NO3)2 (30 ml) x (1.50 M CaCl2) = 45 mmol CaCl2 (15.0 ml) x (0.100 M AgNO3) = 1.50 mmol AgNO3 1.50 millimoles of AgNO3 would react completely with 1.50 x (1/2) = 0.75 millimoles of CaCl2, but there is much more CaCl2 present than that, so CaCl2 is in excess and AgNO3 is the limiting reactant. (0.0150 L) x (0.100 mol/L AgNO3) x (2 mol AgCl / 2 mol AgNO3) x (143.3212 g AgCl/mol) = 0.215 g AgCl